OpenStudy (anonymous):

solve cos(cos^-1X + sin^-1X)

myininaya (myininaya):

remember cos(x+y)=cosxcosy-sinxsiny so we have cos(cos^-1(x))cos(sin^-1(x))-sin(cos^-1(x))sin(sin^-1(x)) =xcos(sin^-1(x))-xsin(cos^-1(x)) ------------------------------- let sin^-1(x)=@ then sin@=x=x/1 so if we draw a right triangle and we label one of the angles @(not the right angle) then opposite side of @ is x and hyp of that triangle is 1 since sin@=opp/hyp so the other can be found using pythagorean thm opp^2+adj^2=hyp^2 \[x^2+adj^2=1^2\] \[adj=\sqrt{1-x^2}\] so cos@=sqrt{1-x^2}/1=sqrt{1-x^2} so \[\cos(\sin^{-1}x)=\cos(@)=\sqrt{1-x^2}\] ----------- now let cos^-1(x)=& so cos&=x=x/1 remember cos&=adj/hyp again assume we have a right traingle we can find the oppsite side of & by using pythagorean thm so we have opp^2+adj^2=hyp^2 \[opp^2+x^2=1^2\] \[opp=\sqrt{1-x^2}\] so \[\sin\&=\sqrt{1-x^2}\] \[\sin(\cos^{-1}x)=\sin(\$)=\sqrt{1-x^2}\] ------------------------- finally \[\cos(\cos^{-1}x+\sin^{-1}x)=\cos(\cos^{-1}x)\cos(\sin^{-1}x)-\sin(\sin^{-1}x)\sin(\cos^{-1}x)\] \[=x*\sqrt{1-x^2}-x*\sqrt{1-x^2}=0\]

OpenStudy (anonymous):

Isn't cos(cos^-1 x) equal to x? (inverse just undoes the operation.

OpenStudy (anonymous):

No that's true but rubbish, I can't split it up like that (time for bed).

OpenStudy (bahrom7893):

estudier are u here.. can u take a look at my lin algebra problem??

OpenStudy (anonymous):

If we let Cos a = x and Sin b = x (ie a = Cos^-1 x, b = Sin^-1 y) then we want Cos (a+b) = Cos^2 a - Sin^2 b = (Cos a + Sin b)(Cos a - Sin b) = 2x * 0 = 0.

myininaya (myininaya):

yes i have cos(cos^(-1)x)=x and sin(sin^(-1)x)=x

OpenStudy (anonymous):

Typo, "b = Sin^-1 y" in my answer should read "b = Sin^-1 x"